OpenStudy (anonymous):

This is a word problem. AIDS cases from 1993 to 2003 the cumulative number N of AIDS cases in thousands can be approximated by N=-2x^2+76x+430, where x=0 corresponds to the year 1993. Year 1993 Cases 422 Year 1995 Cases 565 It goes on but I need help starting!

OpenStudy (anonymous):

Hello is anyone there?

OpenStudy (anonymous):

yes kellyjohnson4, just lookin now!

OpenStudy (anonymous):

okay so, what is the question ?

OpenStudy (anonymous):

for1993 or year zero (x=0), N=-2x^2+76x+430, N=-2(x)^2+76(0)+430 N= -2(0) +76(0) +430 N=0+0+430 N=430 in 1993 or year zero. for 1994 or year one (x=1), N= -2(1)^2+ 76(1)+430 N=-2(1)+76(1)+430 N=-2 +76 +430 N= 74+430 N= 504 For 1995, year 2 (x=2) N= -2(2)^2 +76(2)+430 N=-2(4)+76(2)+430 N=-8+152+430 N=144+430 N=574 etc., these are the first three values using the approximation formula.